1. Find the dimension of a rectangle with perimeter 200 m and whose area is as large as possible



2. The sum of two positive numbers is 10. Find the numbers if the sum of their squares is a minimum

1
Expert's answer
2022-04-11T13:52:41-0400

1. Let the length of the rectangle be x m.


Then width becomes (200 - 2x)/2 = (100 - x) m.


Let the area be y = length times width = x(100-x) sq m.


A plot of y = 100x - x2


shows y (i.e. area) is max when x = 50 m.


Dimensions for maximum area is 50 x 50 meters.


Area is 2,500 m2.

2. If the sum of two positive numbers, x and y is 10


x + y = 10

y = 10 - x


The sum of their squares is"x^2 + (10 - x)^2 = x^2 + 100 - 20x + x^2\n= 2x^2 - 20x + 100"

The minimum will occur when the derivative = 0


f’(x) = 4x - 20

4x - 20 = 0


To get the value of x,


4x = 20

4x/4 = 20/4

x = 5

That is the minimum occurs at (x,y) = (5,5)


and the minimum possible value of the sum of the squares is


52+ 52 = 25 + 25 = 50